OpenStudy (anonymous):

Solve for c and simplify completely. - 10c - 6d = 12

OpenStudy (anonymous):

Add +6d on both sides -10c=12+6d divide by -10 on both sides c = -(12+6d)/10

OpenStudy (anonymous):

\[c = -\frac{ 12+6d }{ 10 }\]

OpenStudy (anonymous):

i believe the answer would be D but not sure

OpenStudy (anonymous):

\[c =- \frac{ 3 }{ 5 }(d+2)\]

OpenStudy (anonymous):

im confushed whered u get the 2 from

OpenStudy (anonymous):

nevermind